PLEASE HELP ME W/ THIS PROBLEM. THANK YOU.
REDUCE 45/75 TO LOWEST TERMS.

Answer by AnlytcPhil(1116)

"PLEASE HELP ME W/ THIS PROBLEM. THANK YOU.
REDUCE 45/75 TO LOWEST TERMS.
`
Think of a whole number smaller than the numerator 45 that
will divide evenly into 45.
`
I thought of 5. So write
`
45 = 5×9
`
There is no integer smaller than 5 that will divide evenly
into 5. So think of a number smaller than 9 that will divide
evenly into 9.
`
I thought of 3. So write
`
45 = 5×9 = 5×3×3
`
Now there are no whole numbers smaller than 3 that will divide
evenly into 3, so we have prime factored 45 as 5×3×3
`
Now lets do the same thing with the denominator 75.
`
Think of a whole number smaller than 75 that will divide evenly
into 75.
`
I thought of 5. So write
`
75 = 5×15
`
There is no integer smaller than 5 that will divide evenly
into 5. So think of a number smaller than 15 that will divide
evenly into 15.
`
I thought of 5. So write
`
75 = 5×5×3
`
Now there are no whole numbers smaller than 3 that will divide
evenly into 3, so we have prime factored 75 as 5×5×3
`
So we have
`
`45 ` ` 5×3×3
---- = --------
`75 ` ` 5×5×3
`
Now we cancel the 5's
`
` ` ` ` 1
`45 ` ` 5×3×3
---- = --------
`75 ` ` 5×5×3
` ` ` ` 1
`
That leaves
`
`45 ` ` 3×3
---- = -----
`75 ` ` 5×3
`
Now cancel the 3's
`
` ` ` ` 1
`45 ` ` 3×3
---- = -----
`75 ` ` 5×3
` ` ` ` ` 1
`
That leaves
`
`45 ` ` 3
---- = ---
`75 ` ` 5
`
So that's the answer, 3/5.
`
Edwin